OpenStudy (anonymous):

Of all positive integers x, for which x is x^2+1 a prime number? (Hint: use induction, starting with 1^2+1=2 is prime.)

OpenStudy (anonymous):

I think I can prove that all of the rest must be even

OpenStudy (anonymous):

not all evens satisfy the condition, though.

OpenStudy (anonymous):

\[(x + n)^2 + 1 = P => x^2 + 2xn + n^2 + 1 = P \]

OpenStudy (anonymous):

\[\frac{P-1}{2} = \frac{x^2 + n^2}{2} + xn\]

OpenStudy (anonymous):

\[\frac{P-1}{2} \] is an integer \[xn\] is an integer \[=> \frac{n^2 +x^2}{2} \] is an integer

OpenStudy (anonymous):

\[=> x^2 + n^2 \] must be even

OpenStudy (anonymous):

=> if x is odd, n is odd and if x is even, n is even

OpenStudy (anonymous):

but once you increment by an odd, you end up at an even

OpenStudy (anonymous):

and when you are at an even, incrementing by an even gives you an even still

OpenStudy (anonymous):

of course, I left out x = 1

OpenStudy (anonymous):

1 is still odd :P